Some hydrogen gas has a volume of 200cm3 at 150c. If the pressure remain constant at what temperature will its volume be 150cm3?
1
Expert's answer
2020-06-05T11:52:20-0400

Given data are

"V_1=200cm^3=2\times 10^{-4}m^3\\T_1=150^{\circ}C=(150+273)=423K\\
V_2=150cm^3=1.5\times 10^{-4}m^3"

Let "T_2" is the temperature of hydrogen gas when volume of gas is "V_2=150cm^3" .



From, The Charles's Law of constant pressure we get

"\frac{V_1}{T_1}=\frac{V_2}{T_2}\\
\implies T_2=\frac{V_2 T_1}{V_1}"

Thus, On plugin the given data we get,

"T_2=\frac{1.5\times 10^{-4}m^3}{2\times 10^{-4}m^3}\times 423\\
\implies T_2=317.25K"